OpenStudy (meowlover17):

Simplify. 3 square root of 10 end root plus 7 square root of 15 end root minus 6 square root of 10 end root minus 4 square root of 15

OpenStudy (meowlover17):

\[3\sqrt{10}+7\sqrt{15}-6\sqrt{10}-4\sqrt{15}\]

OpenStudy (anonymous):

\[3\sqrt{10}+7\sqrt{15}-6\sqrt{10}-4\sqrt{15}\]

OpenStudy (meowlover17):

Yep

OpenStudy (anonymous):

\[=3\sqrt{15}-3\sqrt{10}=3(\sqrt{15}-\sqrt{10})\]

OpenStudy (meowlover17):

So the answer 3 square root of 10 end root minus 3 square root of 15 3 square root of 15 end root minus 3 square root of 10 3 square root of 30 end root minus 3 square root of 20 3 square root of 20 end root minus 3 square root of 30

OpenStudy (meowlover17):

would be answer choice B.

OpenStudy (anonymous):

yes, right

OpenStudy (meowlover17):

Thank you.

OpenStudy (meowlover17):

I'll let you know if i need help later if your online. Appreciate it have a good one :D

OpenStudy (anonymous):

but if u go on it would be " \[3\sqrt{5}(\sqrt{3}-\sqrt{2})\]

OpenStudy (anonymous):

did u become my fan :) ? @MeowLover17

OpenStudy (meowlover17):

Of course

OpenStudy (anonymous):

i didn't take the notification, maybe it will come later ! Thanks anyway ;)

OpenStudy (meowlover17):

Hah, no problem you will see it later, i'm gonna request some answers from you in a little appreciate your help.

OpenStudy (meowlover17):

Actually i need some help with another lesson if you could help that would be great.

OpenStudy (anonymous):

sorry, but i got to go off now ! It's late here in europe now ! :((((

OpenStudy (meowlover17):

All good, have a great night and thanks for your help once again.

OpenStudy (anonymous):

Thanks, you too ! I will help u anytime I'm on! Have a nice day/night ! :P

OpenStudy (meowlover17):

You too.